In the parallelogram, the length AE is 6, the length of BC is 9, and the base DC is 4. Find the area.
Length 6
Base 4
Diagnol length 9
HELP!

Answer by checkley77(12569)

Length 6
Base 4
Diagonal length 9
What you have is a non-right triangle thus the area equation of the triangle is:
area=sqrt[s(s-a)(s-b)(s-c)]
s=(a+b+c)/2
s=(4+6+9)/2=19/2=9.5
Because sis not a whole number the sides must be doubled and then divide the answer by 4.
s=(8+12+18)/2=38/2=19
area=sqrt[19(19-8)(19-12)(19-18)
area=sqrt[19*11*15*7*1]
area=sqrt(21,945)
area=148.14/4=37.03 ans. for half the parallogram.
37.03*2=74.06 ans. for the area of the parallogram.